OpenStudy (anonymous):

Find Amplitude, Period, Phase Shift & Graph one period. y= 1/2 Cos(2x+pi)

OpenStudy (anonymous):

Amplitude= 1/2 Period= pi Keypoints: period/4= \[\frac{ \pi }{ 4 }\] Phase Shift= \[\frac{ \pi }{ 2 }\] How will I find the remaining keypoints & graph!

OpenStudy (primeralph):

what do you mean by keypoints?

OpenStudy (anonymous):

the coordinates on the graph. It has to go by increments of pi/4 for the period pi. If that makes any sense

OpenStudy (primeralph):

the graph is defined for every x, so just plug in an x and get a y. Do you mean specific points?

OpenStudy (anonymous):

x1= pi/2

OpenStudy (anonymous):

x2= pi/2+pi/4

OpenStudy (anonymous):

and so on until you get to x5.. I just want to make sure im not messing up while adding the fractions